OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −x − 6 y = −x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (anonymous):

Type them all in on a graphing calculator and then see if they touch that point

OpenStudy (anonymous):

Ok Ill go find one.

OpenStudy (jim766):

solve for y 3x - y = -3 + y = y 3x = y -3 +3 +3 y = 3x + 3 parallel lines have same slope

OpenStudy (jim766):

so it cant be either of the 1st two sub in the x and y of the point to see which one is true

OpenStudy (anonymous):

Thanks!